OpenStudy (anonymous):

for the given cost function C(x)=16sqrt(x)+(x^2/3375) find a) the cost at the production level 2000. b) the average cost at the production level 2000. c) the marginal cost at the production level 2000. d) the production level that will minimize the average cost. e) the minimal average cost

OpenStudy (cruffo):

\[C(x) = 16\sqrt{x} + \frac{x^2}{3375}\] a.) Find \(C(2000)\). For this part just replace the variable \(x\) with the number \(2000\).

OpenStudy (anonymous):

i tried doing that but it says my answer is wrong

OpenStudy (cruffo):

What are you getting?

OpenStudy (anonymous):

320* sqrt(5) +(3200/27)

OpenStudy (cruffo):

wow...ok:) I would have thought you could just use the calculator for this (application problem). I was thinking the decimal answer of about $1,900.73

OpenStudy (anonymous):

ohh thats right! can you help me with the other parts

OpenStudy (cruffo):

Sure.

OpenStudy (cruffo):

b.) Average cost = \(\overline{C(x)} = \dfrac{C(x)}{x} \) \[= \dfrac{16\sqrt{x} + \dfrac{x^2}{3375}}{x}\] Best to simplify this a bit. \[\dfrac{16\sqrt{x} }{x}+ \dfrac{x^2}{3375x}\] Ok so far?

OpenStudy (anonymous):

i got part b

OpenStudy (anonymous):

yeah i got that answer i need help with c d and e

OpenStudy (cruffo):

Ok.. so I simplified part b to \[\overline{C(x)} = \frac{16}{\sqrt{x}} + \frac{x}{3375}\] c.) Marginal cost = derivative of \(\overline{C(x)}\)

OpenStudy (anonymous):

how can ypu get the derivit for c.) ?

OpenStudy (cruffo):

crap.. marginal cost = derivative of cost marginal average cost = derivative of average cost ...

OpenStudy (cruffo):

\[C(x) = 16\sqrt{x} + \frac{x^2}{3375}\] \[\large C(x) = 16x^{1/2}+ \frac{1}{3375}x^2\] use power rule: \[\large C'(x) = 16\left( \frac{1}{2}x^{-1/2}\right) + \frac{1}{3375}(2x)\]

OpenStudy (anonymous):

so is that the answer to c?

OpenStudy (cruffo):

Simplify a bet and plug in 2000.

OpenStudy (anonymous):

its coming out wrong

OpenStudy (anonymous):

is the 1.36 for part c?

OpenStudy (cruffo):

yes

OpenStudy (anonymous):

yeah its coming out wrong

OpenStudy (anonymous):

but ill skip that for now, do you now how to do part d and e?

OpenStudy (cruffo):

are they maybe asking for marginal average cost???

OpenStudy (anonymous):

the question is asking for marginal cost but ill try entering marginal average cost how do i find that?

OpenStudy (cruffo):

You would find the derivative of average cost using power rule. It's a very small #. Rounding it would be $0.00!! http://www.wolframalpha.com/input/?i=derivative%28+%2816*sqrt%28x%29+%2B+%28x%29^2%2F3375%29%2Fx%2Cx%29+at+x+%3D+2000

OpenStudy (anonymous):

its still coming out wrong :/

OpenStudy (cruffo):

:( not sure then.

OpenStudy (anonymous):

can i do the other parts skipping this part

OpenStudy (cruffo):

d) the production level that will minimize the average cost. This requires taking the derivative of average cost. Then set the derivative equal to 0 and solve for x:

OpenStudy (cruffo):

\[\overline{C(x)} = \frac{16}{\sqrt{x}} + \frac{x}{3375}\] \[\large \overline{C(x)} = 16x^{-1/2}+ \frac{1}{3375}x\] using power rule: \[\large \overline{C'(x)} = 16\left( -\frac{1}{2}x^{-3/2}\right)+ \frac{1}{3375}\] \[\large \overline{C'(x)} = -8x^{-3/2} + \frac{1}{3375}\]

OpenStudy (anonymous):

so we set that equal to 0?

OpenStudy (cruffo):

yep.

OpenStudy (cruffo):

\[ \large -8x^{-3/2} + \frac{1}{3375} = 0\] \[\large \frac{1}{3375} =8x^{-3/2}\] multiply both sides by 1/8 \[\large \frac{1}{27000} =x^{-3/2}\] then raise both sides to the -2/3 power \[\large \left(\frac{1}{27000}\right )^{-2/3} =\left(x^{-3/2}\right) ^{-2/3}\]

OpenStudy (anonymous):

and then the x^(-3/2) cancels out?

OpenStudy (cruffo):

yep, should get x = 900

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

how do i get minimal average cost?

OpenStudy (cruffo):

e) the minimal average cost Replace x with 900 in the average cost function: \[\overline{C(900)} = \frac{16}{\sqrt{900}} + \frac{900}{3375} = $0.80\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

can you help me with a few more questions

OpenStudy (cruffo):

boyfriend just got home...gonna chill for now :) Good luck!

OpenStudy (anonymous):

thanks